# Bifurcations of twisted solutions in a continuum limit for the Kuramoto model on nearest neighbor graphs

### v3

@@ 1. Introduction
Substituting (1.8) into (1.6), we easily see that Eq. (1.8) gives a one-parameter family of solutions, which rotates with the speed $\Omega\neq 0$ when $\sigma\neq 0$ even if $\omega=0$, in the CL (1.6) with (1.7). We take $\kappa$ as a control parameter and prove that the CL (1.6) suffers bifurcations at which the one-parameter family (1.8) of twisted solutions becomes unstable and a stable or unstable two-parameter family of modulated twisted solutions that oscillate or not depending on whether $\sigma\neq 0$ or not are born. It follows from the results of [12, 30, 31] that such bifurcations also occur in the KM (1.1). We demonstrate our theoretical results by numerical simulations for the KM (1.1) on nearest neighbor graphs which may be deterministic dense, random dense or random sparse.
@@ 2. Previous Fundamental Results
    $\min_{\theta\in\mathbb{S}^{1}}|u_{k}^{n}(0)-\bar{u}_{k}^{n}(0)-\theta|<\delta_{1},\quad k\in[n],\quad\mbox{a.s.}$
@@
    then $\bar{\mathbf{u}}(t)$ is stable. Moreover, if for any $\varepsilon>0$, there exists $\delta_{2}>0$ such that for $n>0$ sufficiently large, any solution $u_{k}^{n}(t)$, $k\in[n]$, to the KM (1.1) with
    $\min_{\theta\in\mathbb{S}^{1}}|u_{k}^{n}(t)-\bar{u}_{k}^{n}(t)-\theta|<\delta_{2},\quad k\in[n],\quad\mbox{a.s.}$
    converges to $\bar{u}_{k}^{n}(t)$, $k\in[n]$, then $\bar{\mathbf{u}}(t)$ is asymptotically stable.
@@
    If $\bar{\mathbf{u}}(t)$ is stable, then for any $\varepsilon,T>0$ there exists $\delta>0$ such that for $n>0$ sufficiently large, if $\mathbf{u}_{n}(t)$ is a solution to the KM (1.1) satisfying
    $\min_{\theta\in\mathbb{S}^{1}}\|\mathbf{u}_{n}(0)-\bar{\mathbf{u}}_{n}(0)-\boldsymbol{\theta}\|<\delta\quad\mbox{a.s.},$
@@
    $\min_{\theta\in\mathbb{S}^{1}}\|\mathbf{u}_{n}(t)-\bar{\mathbf{u}}_{n}(t)-\boldsymbol{\theta}\|<\varepsilon\quad\mbox{a.s.}.$
    Moreover, if $\bar{\mathbf{u}}(t)$ is asymptotically stable, then
@@
$\min_{\theta\in\mathbb{S}^{1}}\|\mathbf{u}_{n}(0)-\bar{\mathbf{u}}(0)-\boldsymbol{\theta}\|<\delta\quad\mbox{a.s.},$
@@ 3. Linear Eigenvalue Problem
for some $\ell\in\mathbb{N}$, then $\mathscr{L}$ has a zero eigenvalue of geometric multiplicity three when $\sigma=0$, and a simple zero and a pair of purely imaginary eigenvalues when $\sigma\neq 0$ and $\chi_{2}(\kappa;\ell,q)\neq 0$, so that a bifurcation may occur in the CL (1.6). On the other hand, when $q=0$, the zero eigenvalue of the linear operator $\mathscr{L}$ is always simple and the remaining eigenvalues are negative or positive, depending on $\cos\sigma$ is positive or negative, as shown in Appendix A. Thus, no bifurcation of the solutions (1.8) with $q=0$ occurs in the CL (1.6). Henceforth we only consider the case of $q\in\mathbb{N}$.
@@
where the upper or lower signs are taken simultaneously. Hence, if $c\neq 0$, then $\cos\zeta^{\prime}=1$ so that
@@ 5.1. Case (i): Deterministic graphs
Figure 10. Small oscillations in the steady states of the KM (1.1) with $n=1000$ and $\sigma=\pi/3$ for $q=2$ in case (i): (a) $\kappa=0.16$; (b) $0.166$. They are plotted as the blue and red lines for $k=450$ and $550$, respectively, in Fig. (a) (resp. in Fig. (b)).
@@ 5.2. Cases (ii) and (iii): Random graphs
Figure 13. Time histories of the KM (1.1) with $n=2000$ in case (iii): (a) $(q,\kappa,\sigma)=(1,0.31,0)$; (b) $(2,0.15,0)$; (c) $(1,0.31,\pi/3)$; (d) $(2,0.15,\pi/3)$. They are plotted for every $200$th node (from 100th to 1900th). See also the caption of Fig. 12 for the value of $\omega$.
@@ Appendix A Eigenvalues of $\mathscr{L}$ when $q=0$
for each $\ell\in\mathbb{N}$. Thus, when $\kappa$ is changed, the zero eigenvalue is always simple and the remaining eigenvalues are negative or positive, depending on $\cos\sigma$ is positive or negative.
@@
We substitute (4.2) into (C.1) and integrate the resulting equation with respect to $x$ from 0 to $1$ after multiplying it with $\cos 2\pi j$ or $\sin 2\pi j$, $j\in\mathbb{N}$. Using the above computations, we obtain (4.3) after lengthy calculations.

### v4

@@ 1. Introduction
Substituting (1.8) into (1.6), we easily see that Eq. (1.8) gives a one-parameter family of solutions, which rotates with the speed $\Omega\neq 0$ when $\sigma\neq 0$ even if $\omega=0$, in the CL (1.6) with (1.7). We take $\kappa$ as a control parameter and prove that the CL (1.6) suffers bifurcations at which the one-parameter family (1.8) of twisted solutions becomes unstable and a stable or unstable two-parameter family of modulated twisted solutions that oscillate or not, depending on whether $\sigma\neq 0$ or not are born. It follows from the results of [12, 30, 31] that such bifurcations also occur in the KM (1.1). We demonstrate our theoretical results by numerical simulations for the KM (1.1) on nearest neighbor graphs which may be deterministic dense, random dense or random sparse.
@@ 2. Previous Fundamental Results
    $\min_{\theta\in\mathbb{S}^{1}}|u_{k}^{n}(0)-\bar{u}_{k}^{n}(0)-\theta|<\delta_{1},\quad k\in[n],$
@@
    then $\mathscr{U}$ is stable. Moreover, if for any $\varepsilon>0$, there exists $\delta_{2}>0$ such that for $n>0$ sufficiently large, any solution $u_{k}^{n}(t)$, $k\in[n]$, to the KM (1.1) with
    $\min_{\theta\in\mathbb{S}^{1}}|u_{k}^{n}(0)-\bar{u}_{k}^{n}(0)-\theta|<\delta_{2},\quad k\in[n],\quad\mbox{a.s.}$
    converges to $\mathscr{U}_{n}$, $k\in[n]$, then $\mathscr{U}$ is asymptotically stable.
@@
    If $\mathscr{U}$ is stable, then for any $\varepsilon,T>0$ there exists $\delta>0$ such that for $n>0$ sufficiently large, if $\mathbf{u}_{n}(t)$ is a solution to the KM (1.1) satisfying
    $\min_{\theta\in\mathbb{S}^{1}}\|\mathbf{u}_{n}(0)-\bar{\mathbf{u}}_{n}(0)-\boldsymbol{\theta}\|<\delta,$
@@
    $\min_{\theta\in\mathbb{S}^{1}}\|\mathbf{u}_{n}(t)-\bar{\mathbf{u}}_{n}(t)-\boldsymbol{\theta}\|<\varepsilon\quad\mbox{a.s.}$
    Moreover, if $\mathscr{U}$ is asymptotically stable, then
@@
$\min_{\theta\in\mathbb{S}^{1}}\|\mathbf{u}_{n}(0)-\bar{\mathbf{u}}(0)-\boldsymbol{\theta}\|<\delta,$
@@ 3. Linear Eigenvalue Problem
for some $\ell\in\mathbb{N}$, then $\mathscr{L}$ has a zero eigenvalue of geometric multiplicity three when $\sigma=0$, and a simple zero and a pair of purely imaginary eigenvalues when $\sigma\neq 0$ and $\chi_{2}(\kappa;\ell,q)\neq 0$, so that a bifurcation may occur in the CL (1.6). On the other hand, when $q=0$, the zero eigenvalue of the linear operator $\mathscr{L}$ is always simple and the remaining eigenvalues are negative or positive, depending on whether $\cos\sigma$ is positive or negative, as shown in Appendix A. Thus, no bifurcation of the solutions (1.8) with $q=0$ occurs in the CL (1.6). Henceforth we only consider the case of $q\in\mathbb{N}$.
@@
where the upper or lower signs are taken simultaneously. Hence, if $c\neq 0$, then $\cos\zeta^{\prime}=1$. so that
@@ 5.1. Case (i): Deterministic graphs
Figure 10. Small oscillations in the steady states of the KM (1.1) with $n=1000$ and $\sigma=\pi/3$ for $q=2$ in case (i): (a) $\kappa=0.16$; (b) $0.166$. They are plotted as blue and red lines for $k=450$ and $550$, respectively, in Fig. (a) (resp. in Fig. (b)).
@@ 5.2. Cases (ii) and (iii): Random graphs
Figure 13. Time histories of the KM (1.1) with $n=2000$ in case (iii): (a) $(q,\kappa,\sigma)=(1,0.31,0)$; (b) $(2,0.15,0)$; (c) $(1,0.31,\pi/3)$; (d) $(2,0.15,\pi/3)$. They are plotted for every $200$th node (from 100th to 1900th). See also the captions of Figs. 5 and 12.
@@ Appendix A Eigenvalues of $\mathscr{L}$ when $q=0$
for each $\ell\in\mathbb{N}$. Thus, when $\kappa$ is changed, the zero eigenvalue is always simple and the remaining eigenvalues are negative or positive, depending on whether $\cos\sigma$ is positive or negative.
@@
We substitute (4.2) into (C.1) and integrate the resulting equation with respect to $x$ from 0 to $1$ after multiplying it with $\cos 2\pi j$ or $\sin 2\pi j$, $j\in\mathbb{N}$. So we obtain (4.3) after lengthy calculations.